Let a11 x1 + a12 x2 + a13x3 = b1
a21 x1 + a22 x2 + a23 x3 = b2
a31 x1 + a32 x2 + a33 x3 = b3.
Show that if det (A) 6= 0 where det(A) is the determinant of the coefficient matrix;
then x2 = det(A2)
det(A) where det(A2) is the determinant obtained by replacing the second column of det(A)
by (b1; b2; b3)T :
[
1
Expert's answer
2020-05-08T18:53:05-0400

Let us rewrite the linear system in a matrix form

"\begin{pmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{21} & a_{22} & a_{23} \\
 a_{31} & a_{32} & a_{33} \\
\end{pmatrix}
\cdot \begin{pmatrix}
 x_1 \\
x_2 \\
x_3
\end{pmatrix}
=\begin{pmatrix}
 b_1 \\
 b_2 \\
b_3
\end{pmatrix}" or "\,\, A
\cdot \begin{pmatrix}
 x_1 \\
x_2 \\
x_3
\end{pmatrix}
=\begin{pmatrix}
 b_1 \\
 b_2 \\
b_3
\end{pmatrix}."

We know that "\det (A) \ne 0," so we may write

"\begin{pmatrix}
 x_1 \\
x_2 \\
x_3
\end{pmatrix}
=A^{-1}\cdot\begin{pmatrix}
 b_1 \\
 b_2 \\
b_3
\end{pmatrix} = \dfrac{1}{\det (A)} \begin{pmatrix}
 A_{11} & A_{21} & A_{31} \\
 A_{12} & A_{22} & A_{32} \\
 A_{13} & A_{23} & A_{33} \\
\end{pmatrix}
\cdot \begin{pmatrix}
 b_1 \\
 b_2 \\
b_3
\end{pmatrix} ."

Here "A_{ij}" are the elements of an adjugate matrix (https://en.wikipedia.org/wiki/Adjugate_matrix). Therefore, according to the rules of matrix multiplication we get

"x_2 = \dfrac{1}{\det (A)}\cdot (A_{12}b_1 + A_{22}b_2 + A_{32}b_3)=\\
=\dfrac{1}{\det (A)}\cdot\big(-(a_{21}a_{33}-a_{31}a_{23})b_1 + (a_{11}a_{33}-a_{31}a_{13})b_2 - (a_{11}a_{23}-a_{21}a_{13})b_3\big)." (1)

Now let us compute determinant obtained by replacing the second column of det(A) by (b1; b2; b3)T:

"\det\begin{pmatrix}
 a_{11} & b_1 & a_{13} \\
 a_{21} & b_2 & a_{23} \\
 a_{31} & b_3 & a_{33} \\
\end{pmatrix}
= a_{11}(b_2a_{33} - b_3a_{23}) - b_1(a_{21}a_{33}-a_{31}a_{23}) + a_{13}(a_{21}b_3-a_{31}b_2) = \\
= -b_1(a_{21}a_{33}-a_{31}a_{23}) +b_2(a_{11}a_{33}-a_{13}a_{31}) -b_3(a_{11}a_{23}-a_{13}a_{21})."

If we divide the last expression by "\det (A)" , we'll obtain the expression for "x_2," that we've obtained above (see (1)).

Note that we've obtained a formula from Cramer's rule (see https://en.wikipedia.org/wiki/Cramer's_rule#General_case)